Circular No. 5968

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1994I IN NGC 5194 J. A. Phillips, California Institute of Technology; C. J. Chandler and

M. P. Rupen, National Radio Astronomy Observatory, report: “Using the five-element millimeter interferometer at the Owens Valley Radio Observatory, we have observed SN 1994I at 99 and 218 GHz. We obtained synthesis images of the supernova field at 99 GHz on Apr. 5.2–5.7 and at 218 GHz on Apr. 6.2–6.7 UT. At 99 GHz, we detected an unresolved radio source at the position of the supernova (IAUC 5961, 5963) with flux density 11.0 ± 0.6 mJy (3σ). We did not detect a radio source at 218 GHz; the rms noise in the 218-GHz image was 4 mJy/beam. Our flux scale was established by observations of Neptune and Uranus. The systematic uncertainty in the absolute flux scale is ∼ 15 percent. Comparison with radio measurements (IAUC 5963, 5966) shows that the supernova is already optically thin at millimeter wavelengths; therefore, according to the circumstellar interaction model, the flux density at millimeter wavelengths has already passed maximum and should now be decreasing.”

P. M. Rodriguez-Pascual, R. Monier, and W. Wamsteker, International Ultraviolet Explorer (IUE) Observatory, European Space Agency; C. Fransson, Stockholm Observatory; G. Sonneborn, Goddard Space Flight Center; and R. P. Kirshner, Harvard-Smithsonian Center for Astrophysics, report: “Observations of SN 1994I were made with the IUE on Apr. 3.35, 4.9, and 5.13 UT in the long-wavelength range (200-320 nm). A faint point-source continuum spectrum was detected on top of an extended background, most likely associated with the galaxy M51 itself. Although the detailed flux estimates are complicated by the diffuse background, any increase in flux between the first and third observations is limited to <20 percent. Integrated fluxes in 10-nm bands on Apr. 5.13 (including the diffuse, weak background possibly contributing some 40 percent of the flux) are 5.8 (250 nm) and 6.1 (280 nm) × 10^{−15} erg s^{−1} cm^{−2} A^{−1} (error 15 percent). The spectral appearance is very similar to that seen in SN 1980K a few days before optical maximum, and shows none of the typical metal blanketing below 290 nm (as seen in, e.g., SNe 1983N, 1987A, or 1993J). The spectra show strong interstellar absorption lines (Mn II at 259.37 nm, Fe II at 259.94 nm, and Mg II at 279.55–280.27 nm) redshifted by 0.35 nm, consistent with a possible origin in M51.”

1994 April 8 Daniel W. E. Green